# Thompson Sampling via Fine-Tuning of LLMs

### v1

@@ D.3 FAQ Refinement
We use the Qwen3-1.7B model with the system prompt You are a helpful assistant. followed by the prompt Write an FAQ response to the question ”How do I reset my password?”. /no_think. We adopt a learning rate of 1.0E-7. As kernel features, we extract the first $256$ entries of the Qwen3-Embedding-0.6B model (Zhang et al., 2025) with input set to the concatenation of prompt and response. We normalize the result to the unit sphere and add an additional constant $1$ entry to act as a bias. The reward is modeled as the alignment to the unknown ground-truth response To reset your password, go to the login page and click the “Forgot Password” link. Enter your registered email address, and we’ll send you instructions to create a new password. Make sure to check your spam folder if you don’t see the email within a few minutes., computed through the cosine similarity between the full Qwen3-Embedding-0.6B embeddings of the candidate and the ground-truth.
D.4 Protein Search
@@
We use the ProtGPT2 model (Ferruz et al., 2022) with initial token <||endoftext||> and a learning rate of 1.0E-5. As kernel features, we normalize each $1280$ dimensional token embedding of the generated sequence to the unit sphere, average across the sequence length, normalize the result to the unit sphere, and add an additional constant $1$ entry to act as a bias. The reward is set to the negative thermal instability index (Guruprasad et al., 1990) implemented in the BioPython library (Cock et al., 2009).
D.5 Quantum Circuit Design
@@
We use the Qwen2.5-Coder-1.5B model (Hui et al., 2024) prompted with fill-in-the-middle (Bavarian et al., 2022) to generate a Qiskit circuit (Javadi-Abhari et al., 2024):
@@
We adopt a learning rate of 5.0E-6. As kernel features, we take all $210$ two-qubit Pauli observables, which are efficiently simulated by quantum computers, and concatenate an additional code invalidity bit. If the code invalidity bit is set to 1, all other features are set to zero. Finally, we add a constant $1$ entry to act as a bias. The reward is set to the negative energy of the prepared state according to an unknown Hamiltonian. The Hamiltonian is constructed through the following weighted linear combination of Pauli terms:
@@
D.6 Choice of Hyperparameters
As mentioned in the main text, ToSFiT requires minimal task-specific hyperparameter tuning. Indeed, only the learning rate is tuned for each task, based on one-dimensional grid search with search locations $\{5.0\times 10^{-5},1.0\times 10^{-5},5.0\times 10^{-6},1.0\times 10^{-6},5.0\times 10^{-7},1.0\times 10^{-7},5.0\times 10^{-8}\}$. Burn-in periods of $m\in\{4,8,16\}$ were tried out with $16$ being sufficiently large to ensure stable marginal likelihood maximization, yet small enough to not significantly impact sample efficiency due to delayed optimization. The exploration bonus was sweeped across the set $\{1.0,2.0,4.0,8.0,16.0\}$ on a toy task. A bonus of $4.0$ provided the best trade-off between exploration and exploitation, and generalized to strong performance on the reported experiments. The number of steps of gradient ascent per step of Bayesian optimization, $c$, was kept to $1$ in order to minimize the computational overhead compared to Unguided Generation. For the generation batch size $B$, the values $\{4,8,16,32,64\}$ were tried out. $16$ was the smallest size where GPU utilization became acceptably large during auto-regressive generation. However, empirically $B$ did not significantly impact the simple regret trajectories during Bayesian optimization, suggesting that it could possibly be lowered for alternative models and/or hardware. The baseline Soft Actor Critic requires an additional entropy regularization coefficient as hyperparameter and is swept across $\{0.01,0.1,1.0,10.0\}$ for each experimental setup. We only report on the best configuration, i.e., $0.1$ for FAQ Refinement, $1.0$ for Protein Search, and $10.0$ for Quantum Circuit Design.
@@ Appendix E Proofs
$\gamma^{T}:=\max_{L^{T}}I(Y_{L^{T}};R)\geq 2\underbrace{\tfrac{\ln(1+\sigma_{n}^{\text{-}2})}{4}}_{=:1/C_{\sigma_{n}}}\sum_{t=1}^{T}\sigma_{R_{x}^{t}}^{2}.$
@@
The proof generalizes that of Srinivas et al. (2010) from homoscedastic to heteroscedastic noise. First, note that the expression on the right hand side is well-defined, because $\sigma_{R_{x}^{t}}$ only depends on the observation locations $L^{T}$, but not on the observed value. Next, note that $Y_{L^{T}}|R$ is a multivariate normal with independent components of variance $\sigma_{n}^{2}(x^{1}),\ldots,\sigma_{n}^{2}(x^{T})$, thus
@@
$=H[Y_{L^{T\text{-}1}}]+\tfrac{1}{2}\ln(2\pi e(\sigma^{2}_{n}(x^{T})+\sigma_{R_{x}^{T\text{-}1}}^{2})),$
using that $Y_{x^{T}}|Y_{L^{T\text{-}1}}$ is Gaussian with variance $\sigma^{2}_{n}(x^{T})+\sigma_{R_{x}^{T\text{-}1}}^{2}$. Recursively expanding then results in
$I(Y_{L^{T}};R)=\frac{1}{2}\sum_{t=1}^{T}\ln(1+\sigma^{\text{-}2}_{n}(x^{t})\sigma_{R_{x}^{t\text{-}1}}^{2}).$
Finally, by assumption $\sigma_{R_{x}^{t}}\in[0,1]$, allowing to lower bound each summand
$\sigma_{R_{x}^{t\text{-}1}}^{2}\leq\tfrac{1}{2}\ln(1+\sigma^{\text{-}2}_{n}(x^{t})\sigma_{R_{x}^{t\text{-}1}}^{2})\tfrac{1}{2}\underbrace{\tfrac{4}{\ln(1+\sigma^{\text{-}2}_{n}(x^{t}))}}_{=:C_{\sigma_{n}(x^{t})}}.$
Indeed, the inequality is tight for $\sigma_{R_{x}^{t\text{-}1}}\in\{0,1\}$ and holds in-between due to concavity of
$\sigma_{R_{x}^{t\text{-}1}}^{2}\mapsto\ln(1+\sigma^{\text{-}2}_{n}(x^{t})\sigma_{R_{x}^{t\text{-}1}}^{2})-\sigma_{R_{x}^{t\text{-}1}}^{2}\ln(1+\sigma^{\text{-}2}_{n}(x^{t})).$

### v3

@@ B.4 FAQ Refinement
For ToSFiT, the Actor Critics, Unguided Generation, and Post-Generation TS, we use the Qwen3-1.7B/Qwen3-8B model with the system prompt You are a helpful assistant. followed by the prompt Write an FAQ response to the question "How do I reset my password?". /no_think. We adopt a learning rate of 1.0E-7/2.0E-7. As kernel features, we extract the first $256$ entries of the Qwen3-Embedding-0.6B model (Zhang et al., 2025) with input set to the concatenation of prompt and response. We normalize the result to the unit sphere and add an additional constant $1$ entry to act as a bias. The reward is modeled as the alignment to the unknown ground-truth response To reset your password, go to the login page and click the “Forgot Password” link. Enter your registered email address, and we’ll send you instructions to create a new password. Make sure to check your spam folder if you don’t see the email within a few minutes., computed through the cosine similarity between the full Qwen3-Embedding-0.6B embeddings of the candidate and the ground-truth.
B.5 Protein Search
@@
For ToSFiT, the Actor Critics, Unguided Generation, and Post-Generation TS, we use the ProtGPT2 model (Ferruz et al., 2022) with initial token <||endoftext||> and a learning rate of 1.0E-5. As kernel features, we normalize each $1280$ dimensional token embedding of the generated sequence to the unit sphere, average across the sequence length, normalize the result to the unit sphere, and add an additional constant $1$ entry to act as a bias. The reward is set to the negative thermal instability index (Guruprasad et al., 1990) implemented in the BioPython library (Cock et al., 2009).
B.6 Quantum Circuit Design
@@
For ToSFiT, the Actor Critics, Unguided Generation, and Post-Generation TS, we use the Qwen2.5-Coder-1.5B/Qwen2.5-Coder-7B model (Hui et al., 2024) prompted with fill-in-the-middle (Bavarian et al., 2022) to generate a Qiskit circuit (Javadi-Abhari et al., 2024):
@@
We adopt a learning rate of 5.0E-6/1.0E-5. As kernel features, we take all $210$ two-qubit Pauli observables, which are efficiently simulated by quantum computers, and concatenate an additional code invalidity bit. If the code invalidity bit is set to 1, all other features are set to zero. Finally, we add a constant $1$ entry to act as a bias. The reward is set to the negative energy of the prepared state according to an unknown Hamiltonian. The Hamiltonian is constructed through the following weighted linear combination of Pauli terms:
@@
Appendix C Standardized RLOO Is GRPO
Proposition 3.
Appendix D Constant-Time GP Inference
To ensure efficient runtime in large-scale settings and to integrate with deep neural features $\phi:X\to\mathbb{R}^{d}$, we consider Gaussian processes (GPs) with linear kernels over $\mathbb{R}^{d}$ (Rasmussen and Williams, 2006). As we explain in the following, our implementation requires a constant $\Theta(d^{2})$ memory and compute for conditioning on an observation, inferring posterior mean & variance at a point, and even conducting marginal likelihood maximization. In contrast to a straightforward implementation, it does not increase with the number of past observations.
D.1 Setting
D.2 Closed-Form MLM
Proposition 4.
D.3 Bypassing the Covariance Matrix
@@ E.2 Formal Proofs
$\gamma^{T}:=\max_{L^{T}}I(Y_{L^{T}};R)\geq 2\underbrace{\tfrac{\ln(1+\sigma_{n}^{\text{-}2})}{4}}_{=:1/C_{\sigma_{n}}}\sum_{t=1}^{T}\sigma_{R_{x^{t}}^{t-1}}^{2}.$
@@
The proof generalizes that of Srinivas et al. (2010) from homoscedastic to heteroscedastic noise. First, note that the expression on the right hand side is well-defined, because $\sigma_{R_{x^{t}}^{t-1}}$ only depends on the observation locations $L^{T}$, but not on the observed value. Next, note that $Y_{L^{T}}|R$ is a multivariate normal with independent components of variance $\sigma_{n}^{2}(x^{1}),\ldots,\sigma_{n}^{2}(x^{T})$, thus
@@
$=H[Y_{L^{T\text{-}1}}]+\tfrac{1}{2}\ln(2\pi e(\sigma^{2}_{n}(x^{T})+\sigma_{R_{x^{T}}^{T\text{-}1}}^{2})),$
using that $Y_{x^{T}}|Y_{L^{T\text{-}1}}$ is Gaussian with variance $\sigma^{2}_{n}(x^{T})+\sigma_{R_{x^{T}}^{T\text{-}1}}^{2}$. Recursively expanding then results in
$I(Y_{L^{T}};R)=\frac{1}{2}\sum_{t=1}^{T}\ln(1+\sigma^{\text{-}2}_{n}(x^{t})\sigma_{R_{x^{t}}^{t\text{-}1}}^{2}).$
Finally, by assumption $\sigma_{R_{x^{t}}^{t\text{-}1}}\in[0,1]$, allowing to lower bound each summand
$\sigma_{R_{x^{t}}^{t\text{-}1}}^{2}\leq\tfrac{1}{2}\ln(1+\sigma^{\text{-}2}_{n}(x^{t})\sigma_{R_{x^{t}}^{t\text{-}1}}^{2})\tfrac{1}{2}\underbrace{\tfrac{4}{\ln(1+\sigma^{\text{-}2}_{n}(x^{t}))}}_{=:C_{\sigma_{n}(x^{t})}}.$
Indeed, the inequality is tight for $\sigma_{R_{x^{t}}^{t\text{-}1}}\in\{0,1\}$ and holds in-between due to concavity of
$\sigma_{R_{x^{t}}^{t\text{-}1}}^{2}\mapsto\ln(1+\sigma^{\text{-}2}_{n}(x^{t})\sigma_{R_{x^{t}}^{t\text{-}1}}^{2})-\sigma_{R_{x^{t}}^{t\text{-}1}}^{2}\ln(1+\sigma^{\text{-}2}_{n}(x^{t})).$
@@
Let $Y\sim\mathcal{N}(\nu,\lambda^{2}\Sigma)$ and $\mathds{1}=(1,\ldots,1)^{T}\in\mathbb{R}^{s}$. Then for $y\in\mathbb{R}^{s}$, the (marginal) likelihood $p(Y=y)$ is maximized if